OpenStudy (goldrush18):

Find the limit of:

OpenStudy (goldrush18):

im posting it

OpenStudy (goldrush18):

\[\lim_{x \rightarrow 2} x^3-8 \div x^2-6x+8\]

OpenStudy (goldrush18):

x tends to 2

OpenStudy (goldrush18):

it did not specify it just says find the limit

OpenStudy (anonymous):

if you replace x by 2 and get \(\frac{0}{0}\) then factor and cancel

OpenStudy (goldrush18):

@satellite73 i know but thats not the answer i think i would have to factorize the denominator

OpenStudy (goldrush18):

@matt1888 how do i factor the numerator

OpenStudy (experimentx):

\[ \lim_{x \rightarrow 2} \frac{x^3 - 8}{x^2 - 6x + 8} = \lim_{x \rightarrow 2} \frac{(x - 2)(x^2 + 2x + 4)}{(x-2)(x-4)}\] Cancel out \( (x- 2) \) on both numerator and denominator and put the values of x to get your limit.

OpenStudy (anonymous):

and if you do get \[\frac{0}{0}\] when replacing x by 2 you know how it factors, it must be \[\frac{(x-2)(\text{something})}{(x-2)(\text{something else})}\]

OpenStudy (anonymous):

don't forget, if you know a zero of a polynomial you know how it factors

OpenStudy (goldrush18):

thanks @satellite73 @experimentX and @matt1888 :)

OpenStudy (experimentx):

yw

OpenStudy (anonymous):

in other words it is not a miracle that both numerator and denominator have a common factor of \((x-2)\) they must have a common factor

OpenStudy (anonymous):

also don't forget (because you will see it again) that \[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (goldrush18):

yes i do and i have more questions coming up because i'm preparing for Final exams so i really would appreciate the help from you all :)

OpenStudy (goldrush18):

@satellite73 whenever i see a equation like that in the numerator i keep forgetting that rule

OpenStudy (anonymous):

don't forget for the test! also know that \(a^3+b^3=(a+b)(a^2-ab+b^2)\)

OpenStudy (goldrush18):

@matt1888 im gonna write them down until i know it and btw how was your test?

OpenStudy (goldrush18):

if i could i would have given you all a medal :D

OpenStudy (anonymous):

not to be picky @matt1888 (and byw congratulation on finishing your final) but you cannot "foil it out" because a) it is three terms by two terms and b) you are factoring, not multiplying, so unless you know the answer you will probably not be able to get it (although you could divide is suppose)

OpenStudy (goldrush18):

@matt1888 thats awesome i want to aim high in my finals and go in there with confidence but i have some real issues doing graphs, deferentiation and integration and so forth

OpenStudy (goldrush18):

@matt1888 i will let you know when i look back on the topics these are just past paper questions im going through now

OpenStudy (goldrush18):

Calculus at Pre- University